### v1

@@ Abstract
Generative Flow Networks (GFlowNets, GFNs) are a generative framework for learning unnormalized probability mass functions over discrete spaces. Since their inception, GFlowNets have proven to be useful for learning generative models in applications where the majority of the discrete space is unvisited during training. This has inspired some to hypothesize that GFlowNets, when paired with deep neural networks (DNNs), have favourable generalization properties. In this work, we empirically verify some of the hypothesized mechanisms of generalization of GFlowNets. In particular, we find that the functions that GFlowNets learn to approximate have an implicit underlying structure which facilitate generalization. We also find that GFlowNets are sensitive to being trained offline and off-policy; however, the reward implicitly learned by GFlowNets is robust to changes in the training distribution.
@@ 1 Introduction
Generative Flow Networks (GFlowNets, or GFNs) have emerged as a generative modelling framework for learning unnormalized probability mass functions of discrete objects such as graphs, sequences, or sets Bengio et al. (2021; 2023). They have especially shown promise for a wide range of problems and applications where optimization is conducted over very large combinatorial spaces. To name a few, discrete probabilistic modelling (Zhang et al., 2022), molecular discovery (Bengio et al., 2021; Jain et al., 2023), biological sequence design (Jain et al., 2022), and causal discovery (Deleu et al., 2022; 2023; Atanackovic et al., 2023) are areas where the discrete spaces may be intractably large. In these applications, the model will realistically only visit a small fraction of the overall state space. Therefore, understanding how well GFlowNets model and assign probability mass to the unvisited areas of the state space is a critical step. This is fundamentally a question of generalization.
Generalization of deep neural networks (DNNs) in supervised learning and deep learning has been extensively studied (Leshno et al., 1993; Pascanu et al., 2013; Zhang et al., 2017; 2021; Kawaguchi et al., 2017). Although it is still not entirely understood, many mechanisms and intuitive notions for generalization have emerged as the favored schools of thought (Arpit et al., 2017) and inspire this work. Due to their close relationship with GFlowNets, works studying generalization in Reinforcement Learning also inspire this work (Zhang et al., 2018; Packer et al., 2018; Cobbe et al., 2019; Bengio et al., 2020).
It has been hypothesized that GFlowNets work well because they synergistically leverage the generalization potential of DNNs to assign probability mass in unvisited areas of the state space Bengio et al. (2021), and while the works of  Nica et al. (2022) and Shen et al. (2023) probe at this question, they only do so superficially. Indeed, questions regarding the mechanisms of generalization in GFlowNets have yet to be systematically investigated and wholly understood. With this comes a natural question – how well do GFlowNets generalize to unvisited areas of the state space? Thus, unravelling some intuitive notions on why and how GFlowNets generalize forms the basis of this work.
We center our investigation around three primary hypotheses for generalization in GFlowNets:
@@
    GFlowNets generalize well when their self induced training distributions are nice, i.e. sampling on-policy from $P_{F}(s^{\prime}|s;\theta)$, or proportionally to $R(s)$, is ideal.
@@
    The difficulty for GFlowNets to generalize is modulated more by the complexity of the reward (as perceived by a DNN) than the properties of the distribution induced by the reward.
To investigate these hypotheses, we devise an empirical investigation to disentangle some of the mechanisms that lead to generalization when training GFlowNets. We structure our study on the basis of making simplifying assumptions such that we reduce the complexities/moving parts of training GFlowNets. These “moving parts” may otherwise act as confounding variables when investigating mechanisms. This approach helps us isolate specific factors of variation which may play a role in why GFlowNets generalize.
We describe the details of our experimental setup in §3, and to this end, construct comprehensive benchmark tasks with well defined and tractably computable $p(x;\theta),\forall x\in{\mathcal{X}}$, presented in §2.
We note that our proposed empirical study helps us disentangle some of the mechanisms of generalization in GFlowNets, but does not necessarily determine the true underlying causal order of these mechanisms. Nonetheless, we believe this work can pave a road-map for how to approach testing generalization of GFlowNets and view this as a key step towards understanding important properties for this class of algorithms.
Our main contributions are summarized as follows:
@@
For the sake of clarity, we define some terms that we want to make as least ambiguous as possible in §A.
@@ 1.1 Generative Flow Networks
The so-called balance condition in (1) leads to a variety of learning objectives. In this work we primarily use the Sub-trajectory Balance (specifically SubTB(1)) introduced by Madan et al. (2022), which takes the above conditions, parameterizes the $\log$ flow and logits of policies, taking the squared error over all possible subtrajectories:
@@
We note that trajectory balance (TB) is a special case of the above, where only $n=0$ and $m=T$ are used. We also note from (1) that $P_{F}(s^{\prime}|s)=P_{B}(s|s^{\prime})F(s^{\prime})/F(s)\propto F(s^{\prime})$, and note that $F(s)$ is upper bounded (when $P_{B}$ of all its descendant edges is 1) by the sum of the rewards of all its descendant leaves. For a complete overview of GFlowNets, we refer readers to Bengio et al. (2023).
2 A Benchmark Task for Investigating Generalization Behaviours of GFlowNets
As a first step, we present a novel benchmark environment built on a series of graph-based tasks. We use graphs as the foundation of the benchmark tasks as they are a natural choice for compositional discrete objects, and a wide range of combinatorial problems can be expressed as graph generation. Thus, we use a graph environment to define tasks of varying difficulty. This helps us standardize benchmarks for GFlowNets when investigating generalization and learning $p(x)$ since the environment can be held constant while the reward difficulty is varied. However, for completion, we also conduct experiments on two commonly used environments and tasks in GFlowNet literature: the hypergrid and sequence environments. In the following, we present our graph-based benchmark environment and the corresponding tasks considered in this work.
@@ 2.1 A Small Graph Environment
We setup an environment with all possible graphs of size 7 and less, where the nodes can be one of two colors. This environment has a total of 72296 states, which allows us to compute the true distribution $p(x)$ and the learned $p(x;\theta)$ for $x\in{\mathcal{X}}$ relatively quickly (in the order of 10 seconds on a GPU; see §C.1.2). We denote the set of graphs/states as ${\mathcal{X}}$. We define a number of reward functions that induce a variety of challenges for models.
Reward complexity: We define three different reward functions, which we hope to be of varying difficulty. The hardest function, cliques, requires the model to identify subgraphs in the state which are 4-cliques of at least 3 nodes of the same color. The second function, neighbors, requires the model to verify whether nodes have an even number of neighbors of the opposite color. Finally, the third function, counting, simply requires the model to count the number of nodes of each color in the state. We fully define and show the distribution of $\log R(x)$ of the respective tasks in §C.1. We can observe that the distribution of $\log R(x)$ increases in complexity relative to the hardness of the underlying task. [^1] For example, the relative quantity of high reward states (“sparseness” of rewards over states) and the general distribution of $\log R(x)$ (“discontinuity” of reward distribution) are factors that can drive learning difficulty.
We verify that our intuitive notion of hardness for these tasks translates to graph neural networks (GNNs). To do so we train standard graph attention networks (Veličković et al., 2017) with varying capacity to regress to the (log) reward functions. We use a 90%-10% train-test split, and show the resulting test error in Figure 1. We see that the ordering seems consistent with the difficulty of the task.
Figure 1: Training a GNN on 3 different tasks with models of varying capacity. Most of the variance comes from varying capacity. Dashed lines are for the highest capacity models.
@@
Figure 2: Training a GFlowNet (online and on-policy) on 4 different tasks. While ordering is mostly preserved, apparent difficulty depends on the choice of metric.
@@
Structural generalization: The so-called structural generalization hypothesis posits that function approximators like deep neural networks (DNNs) capture regular patterns in the data (Arpit et al., 2017; Zhang et al., 2021), i.e. structure. This enables models to make accurate predictions on similarly structured inputs not present in the training set. In our investigation, we reify this intuition by introducing a structured test set. We generate a test set randomly (with seeds of course) by picking states with at least 6 nodes (this choice is not arbitrary, see §C.1.3) and adding all their descendants to the test set. We do so iteratively until the test set is of the desired size. We include further details in §C.1.3. Generalizing to a “structured” test set this way poses a more challenging problem than simply selecting test samples i.i.d. from the data distribution (Tossou et al., 2023).
@@
Metrics for Performance: We aim to evaluate how well GFlowNets model and learn the distribution $p(x)$ over all states $x\in{\mathcal{X}}$. To do this, we consider 2 distributional metrics: Jensen-Shannon (JS) divergence and mean absolute error (MAE) between $\log p(x)$ and the learned $\log p(x;\theta)$. We describe details for these metrics in §C.4. We also note that when generating graphs, properly taking into account isomorphic actions is essential to learning the right $p(x;\theta)$ (see Ma et al., 2023, for reference).
@@ 2.2 Difficulty of Tasks is Preserved When Training GFlowNets Online and On-policy
We now verify that training GFlowNets on our proposed graph environment and defined tasks yields the task ranking that we’d expect. We train GFlowNets online with SubTB$(1)$ and a uniform $P_{B}$.
To lower bound the modeling complexity, we add a fourth reward, one where $R(x)=1,\forall x\in{\mathcal{X}}$ [^2] Note that, hypothetically, this may be harder than other rewards. For a constant reward, the model has to learn to put equal probability mass everywhere, which means being able to model the entire state space. In contrast, more complex but “peakier” reward functions may be in some sense easier to “get right” if one cares more about modeling high-reward states, since there are presumably fewer of them. Note that for this specific experiment there are no states hidden from the model, i.e. no test set, since the model is able to sample from its policy and explore the entirety of the state space. We will start using a test set in §4.1. Here, we are only verifying the difficulty of the “usual” GFlowNet setup in the respective tasks.
Figure 2 shows interesting results. The ordering of tasks is more or less conserved, but depends on the metric we use to measure the discrepancy between $p(x)$ and $p(x;\theta)$. Using a constant reward indeed lower bounds other rewards, but not by much. This suggests that there is, unsurprisingly, inherent difficulty in modeling the dynamics of the environment.
We find that there are two main axes of difficulty with respect to the task (aside of course from the scale of the problem, which here is kept constant): (1) how difficult the reward is to model, and (2) how the reward is distributed. Hopefully, Figure 1 and Figure 2 are convincing evidence of (1). We will come back to (2) later in §4.2.
Readers are also hopefully now on board that this benchmark is sufficiently complex, and has interesting characteristics for empirical validation and hypothesis testing purposes. We also consider 2 non-graph environments, a hypergrid environment and sequence environment, with reasonably sized states spaces such that we can tractably compute $p(x;\theta)$ exactly. This allows us to investigate GFlowNet generalization behaviours in contexts other than graphs. Since the graph, hypergrid, and sequence environments are all discrete spaces, we expect general findings to be consistent across all tasks. We use tasks and reward functions used in prior GFlowNet work (Bengio et al., 2021; Malkin et al., 2022; Jain et al., 2023). We describe the setups and experimental details of these environments and tasks in §C.2.

### v2

@@ Abstract
Generative Flow Networks (GFlowNets, GFNs) are a generative framework for learning unnormalized probability mass functions over discrete spaces. Since their inception, GFlowNets have proven to be useful for learning generative models in applications where the majority of the discrete space is unvisited during training. This has inspired some to hypothesize that GFlowNets, when paired with deep neural networks (DNNs), have favorable generalization properties. In this work, we empirically verify some of the hypothesized mechanisms of generalization of GFlowNets. We accomplish this by introducing a novel graph-based benchmark environment where reward difficulty can be easily varied, $p(x)$ can be computed exactly, and an unseen test set can be constructed to quantify generalization performance. Using this graph-based environment, we are able to systematically test the hypothesized mechanisms of generalization of GFlowNets and put forth a set of empirical observations that summarize our findings. In particular, we find (and confirm) that the functions that GFlowNets learn to approximate have an implicit underlying structure which facilitate generalization. Surprisingly —and somewhat contradictory to existing knowledge— we also find that GFlowNets are sensitive to being trained offline and off-policy. However, the reward implicitly learned by GFlowNets is robust to changes in the training distribution.
@@ 1 Introduction
Generative Flow Networks (GFlowNets, or GFNs) have emerged as a generative modeling framework for learning unnormalized probability mass functions of discrete objects such as graphs, sequences, or sets (Bengio et al., 2021; 2023). They have show particular promise for a wide range of problems and applications where optimization is conducted over very large combinatorial spaces. To name a few, discrete probabilistic modeling (Zhang et al., 2022), molecular discovery (Bengio et al., 2021; Jain et al., 2023), biological sequence design (Jain et al., 2022), and causal discovery (Deleu et al., 2022; 2023; Atanackovic et al., 2023) are areas where the discrete spaces may be intractably large. In these applications, the model will realistically only visit a small fraction of the overall state space. Therefore, understanding how well GFlowNets model and assign probability mass to the unvisited areas of the state space is a critical step. This is fundamentally a question of generalization – and is yet to be practically investigated within GFlowNets.
It has been hypothesized that GFlowNets work well because they synergistically leverage the generalization potential of DNNs to assign probability mass in unvisited areas of the state space (Bengio et al., 2021). Indeed, the ability of GFlowNets to learn environment structure plays a critical role in their ability to generalize – a familiar result found in reinforcement learning (RL) (Cobbe et al., 2019; Schrittwieser et al., 2020). Although we have some insight into how GFlowNets work, we know of many failure cases of generalization which exist in RL (Zhang et al., 2018; Packer et al., 2018; Bengio et al., 2020). With the close relationship between GFlowNets and RL (Tiapkin et al., 2024; Deleu et al., 2024; Mohammadpour et al., 2024), we can expect to observe similar problems in GFlowNets when it comes to generalization. Hence, constructing a systematic investigation into generalization within GFlowNets is useful for future algorithmic development, advancement in scientific applications, and our overall understanding.
While the works of Nica et al. (2022) and Shen et al. (2023) probe at the question of generalization within GFlowNets, they only do so superficially. Questions regarding the mechanisms of generalization in GFlowNets have yet to be systematically investigated and wholly understood. Thus, unravelling some intuitive notions on why and how GFlowNets generalize motivates this work. We center our investigation around three primary hypotheses for generalization in GFlowNets:
@@
    GFlowNets generalize well only under a narrow set of distributions, which includes, but is not limited to, sampling from $P_{F}(s^{\prime}|s;\theta)$.
@@
    The difficulty for GFlowNets to generalize is modulated more by the complexity of the reward (functionally, the generalization error of a supervised DNN) than the properties of the distribution induced by the reward; e.g. skewness, temperature.
To investigate these hypotheses, we devise an empirical investigation under 3 experimental settings. (1) Distilling (regressing to) the true flows of the environment. This lets us evaluate generalization on an unseen test set of states and test Hypotheses 23. (2) Measuring memorization gaps when learning the true flows while controlling for data and environment structure. Here we can probe the effect of environment structure on generalization to verify Hypothesis 2. Finally, (3) training GFlowNets offline and off-policy under different training distributions. This lets us explore the effects of offline-off-policy training and changes in the training distribution on generalization; testing Hypothesis 1. These experimental settings help us isolate specific factors of variation which may play a role in why GFlowNets generalize.
To conduct our empirical investigation, we propose a set of new graph-based generation tasks to benchmark the performance of GFlowNets for learning unormalized probability mass functions over discrete spaces. In most existing works, it is common to use the hypergrid and sequence environments to assess performance of GFlowNets in their ability to approximate unormalized probability mass functions. However, the hypergrid and sequence environments are fundamentally limited for assessing generalization performance of GFlowNets as their structure yields tasks that are easy to learn and approximate. Because of this, we introduce new graph-based benchmark environment where we can incrementally increase the difficulty of the reward while also ensuring that we can easily construct a sufficiently challenging test set of unseen (left out) states to quantify generalization performance. This new environment provides us with a means to more robustly test the generalization capabilities of GFlowNets compared to the commonly used benchmark environments (e.g. the hypergrid and sequence environments).
We describe the details of our experimental setups in §3. In §2, we describe our set of comprehensive benchmark tasks with well defined and tractably computable $p(x;\theta),\forall x\in{\mathcal{X}}$ [^1] We note that our proposed empirical study helps us disentangle some of the mechanisms of generalization in GFlowNets, but does not necessarily determine the true underlying causal order of these mechanisms. Nonetheless, we believe this work can pave a road-map for how to approach testing generalization of GFlowNets and view this as a key step towards understanding important properties for this class of algorithms. Our main contributions are summarized as follows:
@@ 1.1 Generative Flow Networks
The so-called balance condition in (1) leads to a variety of learning objectives. In this work we primarily use the Sub-trajectory Balance (specifically SubTB(1)) introduced by Madan et al. (2022) and often considered standard, which takes the above conditions, parameterizes the $\log$ flow and logits of policies, taking the squared error over all possible subtrajectories of $\tau=(s_{0},...,s_{T})$:
@@
We note that trajectory balance (TB) is a special case of the above, where only $n=0$ and $m=T$ are used. We also note that $F(s)$ is upper bounded (when $P_{B}$ of all its descendant edges is 1) by the sum of the rewards of all its descendant leaves. For a complete overview of GFlowNets, we refer readers to Bengio et al. (2023). We define terms that we want to make as least ambiguous as possible in §A.
1.2 Related Work
Generalization in deep learning    Generalization of deep neural networks (DNNs) in supervised learning and deep learning has been extensively studied (Leshno et al., 1993; Pascanu et al., 2013; Zhang et al., 2017; 2021; Kawaguchi et al., 2017). Although it is still not entirely understood, certain mechanisms and intuitive notions for generalization have emerged as the favored schools of thought (Arpit et al., 2017) and inspire this work.
Generalization in reinforcement learning    Due to their close relationship with GFlowNets, works studying generalization in Reinforcement Learning also inspire this work (Zhang et al., 2018; Packer et al., 2018; Cobbe et al., 2019; Bengio et al., 2020). For example, generalization behaviours differ when learning the same function through regression or through temporal credit assignment (Bengio et al., 2020), which in some form GFlowNets make use of.
Generalization in GFlowNets    It is uncontroversial that GFlowNets generalize, to some extent. They learn a distribution that matches the reward in-distribution (Bengio et al., 2021), and on test sets (Malkin et al., 2022), are affected by the choice of parameterization and flow distribution (Shen et al., 2023), and more generally are able to perform much better than random search in large scale intractable problems. That being said, no study so far has tried to zoom in on whether anything is unique about generalization within GFlowNets.
2 A Benchmark Task to Investigate Generalization Behaviours of GFlowNets
As a first step, we present a novel benchmark environment built on a series of graph-based tasks. We use graphs as the foundation of the benchmark tasks as they are a natural choice for compositional discrete objects, and a wide range of combinatorial problems can be expressed as graph generation. We define tasks of varying difficulty on a fixed state space, thus holding the environment constant while the reward difficulty is varied. For completeness, we also conduct experiments on two common benchmark tasks in GFlowNet literature: the hypergrid and sequence tasks.
@@ 2.1 A Small Graph Environment
Table 1: Mean Absolute Error (MAE) when training GFlowNets for different GNN architectures.
Model Constant Counting Neighbors Cliques
GAT 0.08 $\pm$ 0.00 0.14 $\pm$ 0.01 0.30 $\pm$ 0.01 0.32 $\pm$ 0.03
GCN 0.48 $\pm$ 0.01 0.48 $\pm$ 0.01 2.23 $\pm$ 0.19 1.53 $\pm$ 0.27
GIN 0.34 $\pm$ 0.01 0.39 $\pm$ 0.01 1.48 $\pm$ 0.09 0.64 $\pm$ 0.05
Table 2: Jenson-Shannon (JS) divergence when training GFlowNets for different GNN architectures.
Model Constant Counting Neighbors Cliques
GAT 0.002 $\pm$ 0.000 0.002 $\pm$ 0.000 0.005 $\pm$ 0.001 0.008 $\pm$ 0.001
GCN 0.031 $\pm$ 0.001 0.037 $\pm$ 0.001 0.339 $\pm$ 0.017 0.240 $\pm$ 0.049
GIN 0.018 $\pm$ 0.000 0.018 $\pm$ 0.000 0.247 $\pm$ 0.007 0.061 $\pm$ 0.007
@@
Figure 1: (a) Training a GNN on 3 different tasks with models of varying capacity. Most of the variance comes from varying capacity. Dashed lines are for the highest capacity models. (b) Training a GFlowNet (online and on-policy) on 4 different tasks. While ordering is mostly preserved, apparent difficulty depends on the choice of metric.
Consider a space of graphs ${\mathcal{X}}$ and an action space ${\mathcal{A}}\subset{\mathcal{X}}\times{\mathcal{X}}$ of additive graph edits. We then define $P_{F}(\cdot|s):A_{s}\to\mathbb{R}^{+};A_{s}=\{(s,s\textquoteright)|(s,s\textquoteright)\in A\}$. The initial state $s_{0}$ is the empty graph (no nodes and no edges). The action space has 3 actions: (i) add a new node, (ii) add a new edge, or (iii) stop/terminate. When adding nodes, there is a choice between two “colors” of nodes. This defines the state graph $G=({\mathcal{X}},{\mathcal{A}})$ for the graph environment; again note that states are themselves graphs.
We setup an environment with all possible graphs of size 7 and less where the nodes can be one of two colors, for a total of 72296 states. This allows us to compute exactly $p(x;\theta)$ for all $x\in{\mathcal{X}}$ relatively quickly (in the order of 10 seconds on a GPU; see §C.1.2).
Reward complexity: We define three different reward functions, which we hope to be of varying difficulty. The hardest function, cliques, requires the model to identify subgraphs in the state which are 4-cliques of at least 3 nodes of the same color. neighbors, requires the model to verify whether nodes have an even number of neighbors of the opposite color. counting, simply requires the model to count the number of nodes of each color in the state. We fully define and show the distribution of $\log R(x)$ of the respective tasks in §C.1. We can observe that the distribution of $\log R(x)$ increases in complexity relative to the hardness of the underlying task. [^2] For example, the relative quantity of high reward states (“sparseness” of rewards over states) and the general distribution of $\log R(x)$ (“discontinuity” of reward distribution) are factors that can drive learning difficulty.
We verify that our intuitive notion of hardness for these tasks translates to graph neural networks (GNNs). To do so we train standard graph attention networks (Veličković et al., 2017) with varying capacity to regress to the (log) reward functions. We also tried GCNs (Kipf & Welling, 2016) and GINs (Xu et al., 2018), but found GATs to perform significantly better. Because of this, we elected to use GATs for our experiments in this work, as they present the best performance on this graph environment. See Tables 1 and 2 for the ablation over GNN architectures and section D.1 for further details. We use a 90%-10% train-test split, and show the resulting test error in Figure 1(a). We observe that the ordering seems consistent with the difficulty of the task.
@@
Structural generalization: The so-called structural generalization hypothesis posits that function approximators like DNNs capture regular patterns in the data (Arpit et al., 2017; Zhang et al., 2021), i.e. structure. This enables models to make accurate predictions on similarly structured inputs not present in the training set. We reify this intuition by introducing a structured test set, which we generate by picking states with at least 6 nodes and adding all their descendants to the test set, until the test set is of the desired size; see §C.1.3 for details. Generalizing to a “structured” test set this way poses a more challenging problem than simply selecting test samples i.i.d. from the data distribution; this is relevant in domains such as molecular graphs (Tossou et al., 2023), where the choice of split (e.g. based on scaffold) practically matters.
@@
Metrics for Performance: To evaluate how well GFlowNets model the distribution $p(x)$, we consider 2 distributional metrics: Jensen-Shannon (JS) divergence and mean absolute error (MAE) between $\log p(x)$ and the learned $\log p(x;\theta)$; see §C.4 for details. Note that when generating graphs, taking into account isomorphic actions is essential to learning the right $p(x;\theta)$ (see Ma et al., 2023, for reference).
Hypergrid and Sequence Environments: While the hypergrid and sequence environments have been commonly used to sanity-check GFlowNet implementations and methods, we believe them to be too simplistic to leverage the generalization potential of the DNNs, as will be seen in our experimental results.
@@ 2.2 Difficulty of Tasks is Preserved When Training GFlowNets Online and On-policy
We verify that training GFlowNets on our proposed graph environment and defined tasks yields the expected task ranking. We train GFlowNets online with SubTB$(1)$ and a uniform $P_{B}$. To lower bound the modeling complexity, we add a fourth reward, one where $R(x)=1,\forall x\in{\mathcal{X}}$ [^3] Note that, hypothetically, this may be harder than other rewards. For a constant reward, the model has to learn to put equal probability mass everywhere, which means being able to model the entire state space. In contrast, more complex but “peakier” reward functions may be in some sense easier to “get right” if one cares more about modeling high-reward states, since there are presumably fewer of them. Note that for this specific experiment there are no states hidden from the model, i.e. no test set, since the model is able to sample from its policy and explore the entirety of the state space. Results using test sets are presented in §4.1. Here, we only verify the difficulty of the “usual” GFlowNet setup for these tasks.
In Figure 1(b) we observe that the ordering of tasks is more or less conserved, but depends on the metric we use to measure the discrepancy between $p(x)$ and $p(x;\theta)$. Using a constant reward indeed lower bounds other rewards, but not by much. This suggests that there is, unsurprisingly, inherent difficulty in modeling the dynamics of the environment. We find that there are two main axes of difficulty with respect to the task (aside of course from the scale of the problem, which here is kept constant): (1) how difficult the reward is to model, and (2) how the reward is distributed. Hopefully, Figure 1(a) and Figure 1(b) are convincing evidence of (1). We will come back to (2) later in §4.2.
Readers are hopefully now on board that this benchmark is sufficiently complex, and is useful for empirical validation and hypothesis testing purposes. As mentioned, we also consider 2 non-graph environments, a hypergrid environment and sequence environment, with reasonably sized states spaces such that we can tractably compute $p(x;\theta)$ exactly. This allows us to investigate GFlowNet generalization behaviours in contexts other than graphs. We use tasks and reward functions used in prior GFlowNet work (Bengio et al., 2021; Malkin et al., 2022; Jain et al., 2023). See §C.2 for details.